# StereoFoley: Object-Aware Stereo Audio Generation from Video

### v3

@@ Abstract
We present StereoFoley, a video-to-audio generation framework that produces semantically aligned, temporally synchronized, and spatially accurate stereo sound at 48 kHz. While recent generative video-to-audio models achieve strong semantic and temporal fidelity, they largely remain limited to mono or fail to deliver object-aware stereo imaging, constrained by the lack of professionally mixed, spatially accurate video-to-audio datasets. First, we develop and train a base model that generates stereo audio from video, achieving state-of-the-art in both semantic accuracy and synchronization. Next, to overcome dataset limitations, we introduce a synthetic data generation pipeline that combines video analysis, object tracking, and audio synthesis with dynamic panning and distance-based loudness controls, enabling spatially accurate object-aware sound. Finally, we fine-tune the base model on this synthetic dataset, yielding clear object–audio correspondence. Since no established metrics exist, we introduce stereo object-awareness measures and validate it through a human listening study, showing strong correlation with perception. This work establishes the first end-to-end framework for stereo object-aware video-to-audio generation, addressing a critical gap and setting a new benchmark in the field.
@@ 1 Introduction
Recent advances in audio generation [1, 2] have demonstrated the ability to synthesize plausible sounds with different conditioning modalities. Our prior work [3] showed diffusion models can generate spatial audio with user-controlled localization. Video-to-Audio (V2A) generative models have achieved strong semantic alignment and temporal synchronization  [4, 5, 6, 7, 8, 9], yet they largely remain restricted to mono audio. Recent efforts explore stereo rendering, albeit via stereo codecs [10, 11] or post-hoc mono-to-stereo upmixing [12], rather than grounding spatialization in object positions. Other efforts move beyond mono to first-order ambisonics—either extracting 2D videos and explicit camera-direction cues from $360^{\circ}$ videos [13] or directly using panoramic videos [14]. A separate line targets 5.1 surround sound [15], but models scene-level sound fields rather than object-synchronous sound generation. To the best of our knowledge, no prior work directly addresses object-aware stereo sound generation—V2A generation for 2D videos, where sounds emanate from the positions of their corresponding video objects in stereo sound filed. We believe this gap exists partly due to the limitations of widely used datasets for V2A such as VGGSound [16] and AudioSet [17], which lack spatially reliable stereo audio content.
@@ 2.1.1 Encoders
We encode audio using an in-house stereo codec inspired by [19, 2], which maps $x_{\text{audio}}$ into latent representations $z\in\mathbb{R}^{T_{z}\times D_{z}}$, where $T_{z}$ and $D_{z}$ denote the temporal and embedding latent dimensions, respectively. $x_{\text{text}}$ and $x_{\text{audio}}$ are also mapped by CLAP [20, 21] into a shared embedding $c\in\mathbb{R}^{D_{c}}$ used for conditioning.
For $x_{\text{video}}$, following the success demonstrated in MMAudio [9], we employ SyncFormer [22], which produces video embeddings $c_{v}\in\mathbb{R}^{T_{v}\times D_{v}}$, with $T_{v}$ and $D_{v}$ temporal and embedding dimensions. Unlike MMAudio, which augments synchronization through aligned RoPE embeddings [23] and ConvMLPs, we adopt a more simple and direct strategy: in our designed temporal resolution of audio and video latents match $T_{v}=T_{z}$, enabling video and audio embeddings direct integration, achieving strong temporal alignment without additional architectural overhead.
@@ 2.1.2 Generative Base
We employ a Diffusion-Transformer (DiT) [24] as our generative base, composed of cascaded blocks that include self-attention, cross-attention, and gated MLP components, together with layer normalization and skip connections, following a similar architecture to [25]. The model accepts multiple conditioning signals: text and audio embeddings from CLAP, and video embeddings from SyncFormer. Projection layers are used to match the feature dimensionality requirements of each conditioning pathway. CLAP features are provided through cross-attention layers, while video conditioning is directly added to the incoming noisy audio latent. The model is trained using the $v$-objective approach [26], minimizing the Mean Squared Error (MSE) between the ground-truth velocity $v$ and the predicted velocity $v_{\theta}(z_{t},t,\{c,c_{v}\})$ at timestep $t$, conditioned on $c$ and $c_{v}$ given the latent variable $z_{t}$:
@@ 2.2 Synthetic data generation pipeline
For each cluster $c_{i,j}$, generic object-specific sounds are first generated using our in-house text-to-audio (T2A) model conditioned on the LLM-derived descriptions $O_{i,j}$. These preliminary sounds are then temporally aligned with the corresponding cluster using our pre-trained StereoFoley-base model, which takes as inputs the segmented video and the generic audio encoded via CLAP to regenerate synchronized audio $a_{i,j}$ that matches the visual stream. We found this two-stage strategy more stable than directly conditioning V2A on segmented video, which often produced inconsistent or hallucinated outputs. Background ambience is generated separately with the same T2A model, conditioned on $B_{i}$, and mixed without temporal alignment since it does not require precise synchronization.
@@
Given cluster’s horizontal centroid trajectory $\{(\bar{w}_{i,j}[k]\}_{k=1}^{K_{i}}$ and pixel masses $\{m_{i,j}[k]\}_{k=1}^{K_{i}}$ sequences, we spatialize the generated audio $a_{i,j}$ into stereo by mapping motion to panning and mass to loudness, while ensuring temporal stability and object permanence.
To ensure stability and reduce jitter and abrupt jumps between frames, we smooth trajectories and masses data using weighted moving average. Missing detections are filled by linear interpolation between the last and next reliable frames, while off-frame or truncated segments are extrapolated using the average recent velocity. Pixel masses are similarly smoothed, with persistence near screen edges to avoid implausible shrinkage when only part of the object is visible. For off-frame periods, we approximate the mass with the last reliable value (or the first upon re-entry). Finally, horizontal positions and masses are up-sampled from the video frame rate to the audio sampling grid. For a scene $s_{i}$ of duration $T_{i}$, this yields $N_{i}=T_{i}\times f_{s}$ frames. of $\{\bar{w}_{i,j}[n],m_{i,j}[n]\}_{n=0}^{N_{i}-1}$.
@@
We model how distance to an object—approximated by its pixel mass—affects sound using a normalized mass term, with additional attenuation only when the object is off-frame:
@@
where $m_{\max}=\max_{k}(\tilde{m}_{i,j}[k])$ is the maximum observed size of the object in the scene. Thus, the mono audio $a_{i,j}[n]$ of given cluster $c_{i,j}$ is rendered into stereo as:
@@ 3 Experiments and Results
Objective Evaluation (BAS) In-frame Off-frame Combined MMAudio 0.07 0.01 0.08 VGGSound original 0.23 0.20 0.23 StereoFoley-base 0.23 0.21 0.23 StereoFoley-obj 0.33 0.30 0.33 Subjective Evaluation (MOS, 1–5) In-frame only Leaves/enters frame All MMAudio 2.24 2.15 2.19 VGGSound original 2.97 2.97 2.97 StereoFoley-base 3.05 2.82 2.93 StereoFoley-obj 3.54 3.37 3.46
@@ 3.1 Datasets
For our StereoFoley-base experiments, we use VGGSound [16], a widely used public dataset for V2A and Foley sound research, containing approximately 200K video examples. Although the dataset nominally provides stereo audio, our analysis revealed that about $27\%$ of videos are effectively mono, with left and right channels nearly identical (average per-file difference $10^{-5}$). Acknowledging this limitation for stereo learning, we nevertheless used VGGSound as the primary training set for the StereoFoley-base model. Additionally, to also demonstrate our model’s potential when trained on a bigger dataset, we trained our base model on a mix of VGGSound and AudioSet [17], a large-scale dataset of about 2M clips. We filtered AudioSet to remove overlap with VGGSound and excluded clips containing speech and music, making the subset more suitable for Foley sound generation research.
For StereoFoley-obj, we observed that object-aware stereo is underrepresented in VGGSound. Using our synthetic pipeline’s object tracker, we analyzed trajectories and sizes to identify suitable samples. We retained videos with no more than three scenes, at least one of which contained a single trackable object that was either stationary off-center or moved beyond $15\%$ from the screen center. We further excluded cases where objects were too small ($\leq 2\%$ of the frame) or too large ($\geq 60\%$). We found that only about $18\%$ of the dataset satisfied these criteria, and we refer to this subset as VGG-obj. Based on our own inspection of VGGSound’s original audio, we estimate that less than $10\%$ of these clips (roughly $1.8\%$ of VGGSound overall) exhibit clear object-aware stereo spatialization, making VGGSound unsuitable for stereo spatialization learning on its own. For this reason, we generated stereo object-aware audio for the VGG-obj subset, replaced the originals, and used the modified dataset to fine-tune StereoFoley-base into StereoFoley-obj.
@@ 3.2 Experiment Setup and Baselines
We used $T=9.56$ second audio-visual segments from VGGSound, with audio resampled to $f_{s}=48$ kHz stereo. We standardized $x_{\text{video}}$ video streams to $f_{v}=25$ fps at a resolution of $H\times W=224\times 224$ to match SyncFormer input requirements. Our stereo codec encoder produced audio embeddings of $T_{z}\times D_{z}=224\times 256$, and SyncFormer with overlapping window produced video embeddings of $T_{v}\times D_{v}=224\times 768$. Our generative base DiT model has 24 attention layers, 256 channels, and an embedding dimension of 1536, totaling approximately 1.11.1B parameters. For StereoFoley-base, we trained the model from scratch for up to $800$ epochs (about one week) with a global batch size of $512$ on 8×NVIDIA A100 GPUs, using AdamW with a learning rate of $10^{-4}$, weight decay of $10^{-3}$, and $2500$ warmup steps. For stage two, StereoFoley-obj model, we fine-tuned the base model for an additional 150 epochs on a modified VGGSound dataset with replaced VGG-obj videos. During training, we employed classifier-free guidance (CFG) [30] with dropout of $0.1$, and alternated conditioning between audio and text CLAP embeddings, drawing each modality with probability $0.5$. For inference, we used $100$ denoising steps with a CFG scale of $6.0$.
As baselines, we compare our system against MMAudio [9] and Kling-Foley [12].We use the large variant of MMAudio with 44.1,kHz, mono output and $\sim$1.03B parameters. Kling-Foley is stereo systems with $\sim$1.5B parameters, making both models comparable in scale with ours. Being current state of the art in temporal alignment, both systems use similar SyncFormer with RoPE method for synchronization. Despite producing stereo output, Kling-Foley neither discusses object-awareness nor reports any related or stereo-specific metrics. Aslo, as the authors did not release code or checkpoints, we include the results reported in their paper on the VGGSound test set.
@@ 3.3 Evaluation Metrics
For the object-aware stereo, we introduce a bin-alignment score (BAS) to evaluate the alignment between object trajectories and audio spatialization. In videos from VGG-obj test, we compare the normalized horizontal object positions $\bar{w}[k]$ (Sec. 2.2) with the normalized audio energy center of mass positions $\bar{a}[k]$, both quantized into three bins: left, center ($\pm 5\%$), and right. The score is defined as: $BAS=\tfrac{1}{K}\sum_{k=1}^{K}\mathbf{1}[\mathrm{bin}(\bar{w}[k])=\mathrm{bin}(\bar{a}[k])]$, effectively showing a fraction of frames where the two bins match. We report BAS separately for both object in-frame and object off-frame segments, as well as the combined score, reflecting both object awareness and permanence.
We complement the BAS score with a subjective evaluation. We randomly selected 10 videos from VGGS-obj test set: 5 where the object remained on screen throughout, and 5 where the object enters or leaves the frame. We compared four systems: VGGSound original audio, MMAudio, StereoFoley-base, and StereoFoley-obj. A user study was conducted with $180$ participants, who rated how well the sound aligned with the visual object on a 5-point scale (1 = poor, 5 = excellent). As most non-expert raters are unfamiliar with explicitly judging stereo consistency, we added hidden qualification clips with mono sound and left-right flipped sounds, disqualifying raters who rated them “good” or “excellent.”
@@ 3.4 Results
Table 1 compares our models with baselines on the VGGSound test set. Overall, StereoFoley performs on par with both baselines, with small advantages in some metrics and slight deficits in others, all three representing the current state-of-the-art in V2A generation. As noted in Section 3.3, comparisons are imperfect: models differ in output format and Kling-Foley does not report every metric. MMAudio, being mono, likely benefits from mono FD backbones and shows lowest scores, while ours and Kling-Foley yield similar $\text{FD}_{\text{PaSST}}$ score. KL differences are negligible across models, while our model outperforms MMAudio in IS score. We show IB score close to Kling-Foley, both being slightly behind MMAudio. Interestingly, although the differences are marginal, our base model attains the best synchronization score despite using simple video and audio latent addition, unlike RoPE or ConvMLPs in prior works. Additionally, StereoFoley-base achieves a Stereo-Score of 0.21, with the object-aware variant reaching 0.24, confirming stronger stereo separation. Training on larger dataset including AudioSet provides only minor gains to our base model, while StereoFoley-Obj shows slight drops in all metics, likely because of substituting part of VGGSound with synthetic VGG-obj data.
We report stereo object alignment results in Table 2. The upper part shows BAS score on the VGG-obj test subset. As expected, MMAudio, which generates mono audio, shows the weakest alignment. StereoFoley-base performs on par with VGGSound original, effectively learning the stereo characteristics of the dataset, while StereoFoley-obj achieves the highest alignment in both on-frame and off-frame segments, improving the combined score from 0.23 to 0.33. This demonstrates that synthetic data helps learn spatial correspondence between visual objects and audio, highlighting a key finding of our study: stereo awareness can be learned with an end-to-end framework when provided with appropriate data.
The lower half of Table 2 reports Mean Opinion Scores (MOS) from our human study ($N=1,341$ ratings from 131 raters, 49 disqualified). Results are shown separately for in-frame and for cases where objects enter or exit. Consistent with BAS results, MMAudio scored lowest, StereoFoley-base matched VGGSound original audio ratings, and StereoFoley-obj achieved the highest scores. A Kruskal–Wallis test followed by Bonferroni-corrected pairwise comparisons confirmed that StereoFoley-obj significantly outperforms all other systems ($p<0.001$), MMAudio underperforms ($p<0.001$), while StereoFoley-base and VGGSound original are statistically indistinguishable ($p=0.80$). No significant effect was found between in-frame and off-frame conditions ($p>0.05$). These results demonstrate that object-aware training yields perceptually meaningful benefits, and that our BAS metric aligns well with human judgment.

### v4

@@ Abstract
We present StereoFoley, a video-to-audio generation framework that produces semantically aligned, temporally synchronized, and spatially accurate stereo sound at 48 kHz. While recent generative video-to-audio models achieve strong semantic and temporal fidelity, they largely remain limited to mono or fail to deliver object-aware stereo imaging, constrained by the lack of professionally mixed, spatially accurate video-to-audio datasets. First, we develop a base model that generates stereo audio from video, achieving performance on par with state-of-the-art V2A models in both semantic accuracy and synchronization. Next, to overcome dataset limitations, we introduce a synthetic data generation pipeline that combines video analysis, object tracking, and audio synthesis with dynamic panning and distance-based loudness controls, enabling spatially accurate object-aware sound. Finally, we fine-tune the base model on this synthetic dataset, yielding clear object–audio correspondence. Since no established metrics exist, we introduce a stereo object-awareness metric and report it alongside a human listening study; the two evaluations exhibit consistent trends. This work establishes the first end-to-end framework for stereo object-aware video-to-audio generation, addressing a critical gap in the field.
@@ 1 Introduction
Recent advances in audio generation [19, 7] have demonstrated the ability to synthesize plausible sounds with different conditioning modalities. Our prior work, ImmerseDiffusion [10], we showed diffusion models can generate spatial audio with user-controlled localization. Video-to-Audio (V2A) generative models have achieved strong semantic alignment and temporal synchronization  [13, 30, 33, 32, 28, 2], yet they largely remain restricted to mono audio. Recent efforts explore stereo rendering, albeit via stereo codecs [27, 21] or post-hoc mono-to-stereo upmixing [29], rather than grounding spatialization in object positions. Other efforts move beyond mono to first-order ambisonics—either extracting 2D videos and explicit camera-direction cues from $360^{\circ}$ videos [15] or directly using panoramic videos [20]. A separate line targets 5.1 surround sound [5], but models scene-level sound fields rather than object-synchronous sound generation. To the best of our knowledge, no prior work directly addresses object-aware stereo sound generation—V2A generation for 2D videos, where sounds emanate from the positions of their corresponding video objects in stereo sound field. We believe this gap exists partly due to the limitations of widely used datasets for V2A such as VGGSound [1] and AudioSet [9], which lack spatially reliable stereo audio content.
@@ 2.1.1 Encoders
We encode audio using an in-house stereo codec inspired by [18, 7], which maps $x_{\text{audio}}$ into latent representations $z\in\mathbb{R}^{T_{z}\times D_{z}}$, where $T_{z}$ and $D_{z}$ denote the temporal and embedding latent dimensions, respectively. $x_{\text{text}}$ and $x_{\text{audio}}$ are mapped by in-house CLAP [31, 6] model into a shared embedding $c\in\mathbb{R}^{D_{c}}$ used for conditioning.
For $x_{\text{video}}$, following the success demonstrated in MMAudio [2], we employ Synchformer [14], which produces video embeddings $c_{v}\in\mathbb{R}^{T_{v}\times D_{v}}$, with $T_{v}$ and $D_{v}$ temporal and embedding dimensions. Unlike MMAudio, which augments synchronization through aligned RoPE embeddings [26] and ConvMLPs, we adopt a more simple and direct strategy: in our designed temporal resolution of audio and video latents match $T_{v}=T_{z}$, enabling video and audio embeddings direct integration, achieving strong temporal alignment without additional architectural overhead.
@@ 2.1.2 Generative Base
We employ a Diffusion-Transformer (DiT) [22] as our generative base, composed of cascaded blocks that include self-attention, cross-attention, and gated MLP components, together with layer normalization and skip connections, following a similar architecture to [8]. The model accepts multiple conditioning signals: text and audio embeddings from CLAP, and video embeddings from Synchformer. Projection layers are used to match the feature dimensionality requirements of each conditioning pathway. CLAP features are provided through cross-attention layers, while video conditioning is directly added to the incoming noisy audio latent. The model is trained using the $v$-objective approach [25], minimizing the Mean Squared Error (MSE) between the ground-truth velocity $v$ and the predicted velocity $v_{\theta}(z_{t},t,\{c,c_{v}\})$ at timestep $t$, conditioned on $c$ and $c_{v}$ given the latent variable $z_{t}$:
@@ 2.2 Synthetic data generation pipeline
For each cluster $c_{i,j}$, generic object-specific sounds are first generated using our in-house text-to-audio (T2A) model conditioned on the LLM-derived descriptions $O_{i,j}$. These preliminary sounds are then temporally aligned with the corresponding cluster using our pre-trained StereoFoley-base model, which takes as inputs the segmented video and the generic audio encoded via CLAP to regenerate synchronized audio $a_{i,j}$ that matches the visual stream. We found this two-stage strategy more stable than directly conditioning V2A on segmented video, which often produced inconsistent or hallucinated outputs. Background (non-object) sounds are generated separately using the same T2A model conditioned on $B_{i}$ and mixed without temporal alignment, since they are not tied to specific visual events.
@@
Given cluster’s horizontal centroid trajectory $\{(\bar{w}_{i,j}[k]\}_{k=1}^{K_{i}}$ and pixel masses $\{m_{i,j}[k]\}_{k=1}^{K_{i}}$, we spatialize the generated audio $a_{i,j}$ using standard cinematic stereo conventions for 2D video, mapping horizontal motion to audio panning and mass to loudness.
To ensure stability and reduce jitter and abrupt jumps between frames, we smooth trajectories and masses data using weighted moving average. Missing detections are filled by linear interpolation between the last and next reliable frames, while off-screen or truncated segments are extrapolated using the average recent velocity. Pixel masses are similarly smoothed, with persistence near screen edges to avoid implausible shrinkage when only part of the object is visible. For off-screen periods, we approximate the mass with the last reliable value (or the first upon re-entry). Finally, horizontal positions and masses are up-sampled from the video frame rate to the audio sampling grid. For a scene $s_{i}$ of duration $T_{i}$, this yields $N_{i}=T_{i}\times f_{s}$ frames of $\{\bar{w}_{i,j}[n],m_{i,j}[n]\}_{n=0}^{N_{i}-1}$.
@@
We model how distance to an object—approximated by its pixel mass—affects sound using a normalized mass term, with additional attenuation only when the object is off-screen:
@@
where $m_{\max}=\max_{k}(\tilde{m}_{i,j}[n])$ is the maximum observed size of the object in the scene. Thus, the mono audio $a_{i,j}[n]$ of given cluster $c_{i,j}$ is rendered into stereo as:
@@ 3 Experiments and Results
Objective Evaluation (BAS) on-screen off-screen Combined MMAudio 0.07 0.01 0.08 VGGSound original 0.23 0.20 0.23 StereoFoley-base 0.23 0.21 0.23 StereoFoley-obj 0.33 0.30 0.33 Subjective Evaluation (MOS, 1–5) on-screen only on/off-screen All MMAudio 2.24 2.15 2.19 VGGSound original 2.97 2.97 2.97 StereoFoley-base 3.05 2.82 2.93 StereoFoley-obj 3.54 3.37 3.46
@@ 3.1 Datasets
For our StereoFoley-base experiments, we use VGGSound [1], a widely used public dataset for V2A and Foley sound research, containing approximately 200K video examples. Although the dataset nominally provides stereo audio, our analysis revealed that about $27\%$ of videos are effectively mono, with left and right channels nearly identical (average channel difference $10^{-5}$). Acknowledging this limitation for stereo learning, we nevertheless used VGGSound as the primary training set for the StereoFoley-base model. Additionally, to also demonstrate our model’s potential when trained on a bigger dataset, we trained our base model on a mix of VGGSound and AudioSet [9], a large-scale dataset of about 2M clips. We filtered AudioSet to remove overlap with VGGSound and excluded clips containing speech and music, making the subset more suitable for Foley sound generation research.
For StereoFoley-obj, we observed that object-aware stereo is underrepresented in VGGSound. Using our synthetic pipeline’s object tracker, we analyzed trajectories and sizes to identify suitable samples. We retained videos with no more than three scenes, at least one of which contained a single trackable object that was either stationary off-center or moved beyond $15\%$ from the screen center. We further excluded cases where objects were too small ($\leq 2\%$ of the frame) or too large ($\geq 60\%$). We found that only about $18\%$ of the dataset satisfied these criteria, and we refer to this subset as VGG-obj. Based on our own inspection of VGGSound’s original audio, we estimate that less than $10\%$ of these clips (roughly $1.8\%$ of VGGSound overall) exhibit clear object-aware stereo spatialization, making VGGSound unsuitable for stereo spatialization learning on its own. We generated stereo object-aware audio for the VGG-obj subset. Replacing only this subset, we fine-tuned StereoFoley-base into StereoFoley-obj while retaining natural audio ambience.
@@ 3.2 Experiment Setup and Baselines
We used $T=9.56$ second audio-visual segments from VGGSound, with audio resampled to $f_{s}=48$ kHz stereo. We standardized $x_{\text{video}}$ video streams to $f_{v}=25$ fps at a resolution of $H\times W=224\times 224$ to match Synchformer input requirements. Our stereo codec encoder produced audio embeddings of $T_{z}\times D_{z}=224\times 256$, and Synchformer with overlapping window produced video embeddings of $T_{v}\times D_{v}=224\times 768$. Our generative base DiT model has 24 attention layers, 256 channels, and an embedding dimension of 1536, totaling approximately 1.11.1B parameters. For StereoFoley-base, we trained the model from scratch for up to $800$ epochs (about one week) with a global batch size of $512$ on 8×NVIDIA A100 GPUs, using AdamW with a learning rate of $10^{-4}$, weight decay of $10^{-3}$, and $2500$ warmup steps. For stage two, StereoFoley-obj model, we fine-tuned the base model for an additional 150 epochs on a modified VGGSound dataset with replaced VGG-obj videos. During training, we employed classifier-free guidance (CFG) [11] with dropout of $0.1$, and alternated conditioning between audio and text CLAP embeddings, drawing each modality with probability $0.5$. For inference, we used $100$ denoising steps with a CFG scale of $6.0$.
As baselines, we compare our system against MMAudio [2] and Kling-Foley [29].We use the large variant of MMAudio with 44.1,kHz, mono output and $\sim$1.03B parameters. Kling-Foley is stereo systems with $\sim$1.5B parameters, making both models comparable in scale with ours. Being current state of the art in temporal alignment, both systems use similar Synchformer with RoPE method for synchronization. Despite producing stereo output, Kling-Foley neither discusses object-awareness nor reports any related or stereo-specific metrics. Also, as the authors did not release code or checkpoints, we include the results reported in their paper on the VGGSound test set.
@@ 3.3 Evaluation Metrics
As no existing spatial AV alignment metrics are directly applicable to open-domain stereo object correspondence, we introduce a bin-alignment score (BAS) to evaluate the alignment between object trajectories and audio spatialization. In videos from VGG-obj test, we compare the normalized horizontal object positions $\bar{w}[k]$ (Sec. 2.2) with the normalized audio energy center of mass positions $\bar{a}[k]$, both quantized into three bins: left, center ($\pm 5\%$), and right. The score is defined as: $BAS=\tfrac{1}{K}\sum_{k=1}^{K}\mathbf{1}[\mathrm{bin}(\bar{w}[k])=\mathrm{bin}(\bar{a}[k])]$, effectively showing a fraction of frames where the two bins match. We report BAS separately for both object on-screen and object off-screen segments, as well as the combined score, reflecting both object awareness and permanence.
We complement the BAS score with a subjective evaluation. We randomly selected 10 videos from VGGS-obj test set: 5 where the object remained on-screen throughout, and 5 where the object enters or leaves the screen. We compared four systems: VGGSound ground truth audio, MMAudio, StereoFoley-base, and StereoFoley-obj. A user study was conducted with $180$ participants, who rated how well the sound aligned with the visual object on a 5-point scale (1 = poor, 5 = excellent). As most non-expert raters are unfamiliar with explicitly judging stereo consistency, we added hidden qualification clips with mono sound and left-right flipped sounds, disqualifying raters who rated them “good” or “excellent.”
@@ 3.4 Results
Table 1 compares our models with baselines on the VGGSound test set. Overall, StereoFoley performs on par with both baselines, with small advantages in some metrics and slight deficits in others, all three representing the current state-of-the-art in V2A generation. As noted in Section 3.3, comparisons are imperfect: models differ in output format and Kling-Foley does not report every metric. MMAudio, being mono, likely benefits from mono FD backbones and shows the least distances, while ours and Kling-Foley yield similar $\text{FD}_{\text{PaSST}}$. KL differences are negligible across models, while our model outperforms MMAudio in IS score. We show IB score close to Kling-Foley, both being slightly behind MMAudio. Interestingly, although the differences are marginal, our base model attains the best synchronization score despite using simple video and audio latent addition, unlike RoPE or ConvMLPs in prior works. Additionally, StereoFoley-base achieves a Stereo-Score of 0.21, with the object-aware variant reaching 0.24, confirming stronger stereo separation. Training on larger dataset including AudioSet provides only minor gains to our base model, while StereoFoley-Obj shows slight drops in all metics, likely because of substituting part of VGGSound with synthetic VGG-obj data.
We report stereo object alignment results in Table 2. The upper part shows BAS score on the VGG-obj test subset. As expected, MMAudio, which generates mono audio, shows the weakest alignment. StereoFoley-base performs on par with VGGSound original, effectively learning the stereo characteristics of the dataset, while StereoFoley-obj achieves the highest alignment in both on-screen and off-screen segments, improving the combined score from 0.23 to 0.33. This demonstrates that synthetic data helps learn spatial correspondence between visual objects and audio, highlighting a key finding of our study: stereo awareness can be learned with an end-to-end framework when provided with appropriate data.
The lower half of Table 2 reports Mean Opinion Scores (MOS) from our human study ($N=1,341$ ratings from 131 raters, 49 disqualified). Results are shown separately for on-screen only and for cases where object enters or exits screen. Consistent with BAS results, MMAudio scored lowest, StereoFoley-base matched VGGSound original audio ratings, and StereoFoley-obj achieved the highest scores. A Kruskal–Wallis test followed by Bonferroni-corrected pairwise comparisons confirmed that StereoFoley-obj significantly outperforms all other systems ($p<0.001$), MMAudio underperforms ($p<0.001$), while StereoFoley-base and VGGSound original are statistically indistinguishable ($p=0.80$). No significant effect was found between on-screen only and on/off-screen conditions ($p>0.05$). These results demonstrate that object-aware training yields perceptually meaningful benefits, and that our BAS metric aligns well with human judgment.